OpenStudy (alexwee123):

f(x)= ln( 2^x x^2) find the derivative

OpenStudy (anonymous):

\(f(x)=\ln(2^x\cdot x^2)\), right? You just use chain rule followed by product rule, and you can rewrite the \(2^x\) as \(e^{x\ln 2}\).

OpenStudy (alexwee123):

ohhh the natural log part just confused me :) thx :D

OpenStudy (anonymous):

No worries. Actually I just realized, to make it easier, you can simplify it first using logarithm rules. That way you don't have to use product rule at all. You can end up with just \(x\ln2+2\ln x\), I think.

OpenStudy (anonymous):

Which is very easy to derive :)

OpenStudy (alexwee123):

haha ya thx XD but my teacher wants me to practice on the product rule cuz i ll be using it a lot more later

OpenStudy (anonymous):

Yes, I agree with your teacher, practicing the product rule is a good idea. Also practice the quotient and chain rules. You should know them like the back of your hand :)

OpenStudy (alexwee123):

aight thx for ur help :)